A block of mass M is pulled along a horizontal frictionless surface by a rope of mass m. If a force
F is applied at one end of the rope, the force which the rope exerts on the block is:

Grade:

1 Answers

Rinkoo Gupta
9 years ago
the pull of the rope is the only force acting on the block. It is also accelerating with the block:

F=M·a=M·Fm+M


Thanks & Regards
Rinkoo Gupta
AskIITians Faculty